• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Jack (Techs)

    Log entry time 18:27:53 on November 18, 2008

    Entry number 249590

    keyword=Right Q2 LHe level meter

    Looking at the installed meter with a scope it clearly had issues on the voltage output stage, 60 Hz ripple. We swapped it out. The new meter was stable on the output and gave us a number we expected. In the true spirit of experimental physics we quit when we got the number we wanted to see. Cryo set the PID loop and we will keep an eye on things.